OpenStudy (anonymous):

Please Help!! Greatly appreciated What sine function represents an amplitude of 1, a period of 2pi, a horizontal shift of pii, and a vertical shift of -4?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hello Ms Chrissy! Our sine function looks like this:\[\Large\rm y=A \sin\left[b(x-c)\right]+d\]Where: \(\Large\rm A\) is the simplitude, \(\Large\rm \frac{2\pi}{b}\) is the period, \(\Large\rm c\) is the horizontal shift, and \(\Large\rm d\) is the vertical shift.

zepdrix (zepdrix):

So ummmm

zepdrix (zepdrix):

Using that basic function, do you think you can try to plug the numbers into the correct spots? :d

OpenStudy (anonymous):

Yes, y= sin (2pi/b(x-pi))-4

zepdrix (zepdrix):

Oh i said that kind of confusing.. sorry sorry.\[\Large\rm period = \frac{2\pi}{b}\]So moving stuff around gives us this equation:\[\Large\rm b=\frac{2\pi}{period}\]And our period was given to be 2pi, yes? So we'll plug that in, which will help us determine our b value,\[\Large\rm b=\frac{2\pi}{2\pi}\]

zepdrix (zepdrix):

You determined that your A value is simply 1, good. Looks like our b value is just 1 also, yah? And your shifts look correct! :)\[\Large\rm y=A \sin\left[b(x-c)\right]+d\]\[\Large\rm y=1 \sin\left[1(x-\pi)\right]-4\]\[\Large\rm y=\sin(x-\pi)-4\]

OpenStudy (anonymous):

Thank you so much for helping me understand, I didn't know it was that simple! Can you help me with another homework problem?

zepdrix (zepdrix):

sure c: let's try another

OpenStudy (anonymous):

Okay, What is the rate of change from x = 0 to x = pi over 2?

zepdrix (zepdrix):

Rate of change between these two points is going to be the slope of the line that passes through them.

zepdrix (zepdrix):

Do you remember your slope formula?

OpenStudy (anonymous):

Okay and yes slope = y2 - y1/x2-x1

zepdrix (zepdrix):

\[\Large\rm m=\frac{y_2-y_1}{x_2-x_1}\] Ok good. Well we know the x values that need to use, yes? Those shouldn't be hard to plug in.\[\Large\rm m=\frac{y_2-y_1}{\frac{\pi}{2}-0}\]What are the y coordinates though? Can you correctly determine the y coordinate at pi/2 by looking at the graph?

OpenStudy (anonymous):

Y=3

OpenStudy (anonymous):

y=0

zepdrix (zepdrix):

\[\Large\rm m=\frac{3-0}{\frac{\pi}{2}-0}=\frac{3}{\pi/2}\]Mmmm ok good good. Do you understand how to simply this expression? Dividing by a fraction can be a little tricky. Did you learn "keep change flip" or something similar maybe?

OpenStudy (anonymous):

I'm not sure. I know I keep pi

zepdrix (zepdrix):